OpenStudy (megannicole51):

solve the integral ((x)/(x^2-9x+20))dx I solved it all the way to the bottom and now i am at x(A+B)+A(-4)+B(-5)=x I was told that A+B can only equal 0 when the number after the equal sign (in this case "x" cannot be a variable). So i tried solving it out anyways and got A=x and B=-x. Then I plugged it into the original ((A)/(x-5))+((B)/(x-4)) and it doesn't work. @pgpilot326

OpenStudy (anonymous):

is this partial fractions?

OpenStudy (anonymous):

\[x^2-9x+20=(x-4)(x-5)\] \[\frac{x}{(x-4)(x-5)}=\frac{A}{x-4}+\frac{b}{x-5}\] there is a real quick snappy way to solve

OpenStudy (anonymous):

that is what you are looking for, right?

OpenStudy (anonymous):

A and B are numbers, not variables. one way to find A is this take \[\frac{x}{(x-4)(x-5)}\] put your hand over the part of the denominator that is the denominator for A \[\frac{x}{\cancel{(x-4)}(x-5)}\] replace \(x\) by \(4\) and get \[\frac{4}{4-5}=-4\]

OpenStudy (megannicole51):

yes it is....sorry my laptop died

OpenStudy (anonymous):

ok so now we have \[\frac{-4}{x-4}\] for the A part

OpenStudy (anonymous):

\[ \frac{x}{x^{2}-9x+20}=\frac{A}{x-5}+\frac{B}{x-4} \Rightarrow A(x-4)+B(x-5)=x\]

OpenStudy (megannicole51):

yup ive got that....then i have x(A+B)+A(-4)+B(-5)=x

OpenStudy (anonymous):

for B repeat and get \[\frac{x}{(x-4)\cancel{(x-5)}}\]

OpenStudy (anonymous):

you get 2 equations in 2 unknowns... this is a longer way as it should yield the same thing @satellite73 was saying.

OpenStudy (anonymous):

put in \(x=5\) to get \[\frac{5}{5-4}=5\]

OpenStudy (megannicole51):

@satellite73 i dont understand why you are crossing them out of the denominator

OpenStudy (anonymous):

ok lets go back to the beginning i was just showing off a snappy trick

OpenStudy (megannicole51):

so @pgpilot326 what do i do after..x(A+B)+A(-4)+B(-5)=x

OpenStudy (anonymous):

\[A(x-5)+B(x-4)=x\] replace \(x\) by \(5\)

OpenStudy (anonymous):

@satellite73 you show off...

OpenStudy (anonymous):

you get \[B(5-4)=5\] sp \[B=5\]

OpenStudy (megannicole51):

i dont need to go to the beginning i just need to know what to do after x(A+B)+A(-4)+B(-5)=x please:)

OpenStudy (anonymous):

that is not correct though

OpenStudy (anonymous):

like i said... you get 2 equations in 2 unknowns... A + B = 1 -4A-5B=0 solve the system

OpenStudy (anonymous):

\[\frac{x}{x^{2}-9x+20}=\frac{A}{x-5}+\frac{B}{x-4}\] so \[A(x-4)+B(x-5)=x\] for the numerator now this has to be true for all \(x\) in particular it must be true of \(x=5\)

OpenStudy (megannicole51):

why is A+B=1?

OpenStudy (anonymous):

Ax+Bx=x so this is the same as A+B=1

OpenStudy (anonymous):

i will shut up except to say this is a long and hard way to do it

OpenStudy (anonymous):

i know

OpenStudy (anonymous):

would you like to do it the easy way? without solving a system of equations?

OpenStudy (anonymous):

the main thing is that @megannicole51 understand and can do it herself... it doesn't matter to me

OpenStudy (megannicole51):

i just need to know how to do it after x(A+B)+A(-4)+B(-5)=x.....ive been doing them like this all day so i dont want to get confused

OpenStudy (anonymous):

ok then as @pgpilot326 said, this means \(A+B=1\) since \((A+B)x=1x\) and also that \(-4A-5B=0\) since there is no constant on the right hand side

OpenStudy (megannicole51):

so if there is no constant A+B will always equal 1 by default?

OpenStudy (anonymous):

x terms with x terms and constant terms with constant terms...

OpenStudy (anonymous):

no it is because \(x=(A+B)x\)

OpenStudy (anonymous):

if you had for example \(3x\) on the right, then \(A+B=3\)

OpenStudy (megannicole51):

oooooh that makes sense

OpenStudy (megannicole51):

okay let me see if i can finish it by myself

OpenStudy (anonymous):

ok, then if you like, i can show you how to do it in your head without writing anything down

OpenStudy (megannicole51):

let me do it real fast then okay!

OpenStudy (anonymous):

k

OpenStudy (anonymous):

you're in good hands @megannicole51 , I'll leave you be

OpenStudy (megannicole51):

thank you for your help ill message you if i need help! :D

OpenStudy (megannicole51):

so why does -4A-5B=0?

OpenStudy (anonymous):

because A and B are numbers so \(-4 A-5B\) is a number (constant) but there is no constant in the numerator, only \(x\)

OpenStudy (megannicole51):

oh okay

OpenStudy (anonymous):

if the numerator was \(x+7\) then \(-4A-5B=7\)

OpenStudy (megannicole51):

gotcha!

OpenStudy (megannicole51):

okay so in order to solve for A and B i just do simple algebra and substitute correct?

OpenStudy (anonymous):

you solve the system however you like elimination or in this case substitution which will probably be easier

OpenStudy (megannicole51):

so i have -4A+5(4A/5)=0....am i on the right track?

OpenStudy (megannicole51):

im doing substitution

OpenStudy (anonymous):

no i don't think so \[A+B=1\iff B=1-A\] substitute in to \[-4A-5B=0\] and solve for \(A\)

OpenStudy (anonymous):

clear or no?

OpenStudy (megannicole51):

yeah it is hold on

OpenStudy (megannicole51):

im having pencil issues lol

OpenStudy (anonymous):

k

OpenStudy (megannicole51):

A=5 B=-4

OpenStudy (anonymous):

whew!!

OpenStudy (megannicole51):

yay!

OpenStudy (anonymous):

you see how long that took? all that work? multiply out, solve a system yechh

OpenStudy (megannicole51):

well i have to do it for tests and quizzes and homework :/

OpenStudy (anonymous):

then lets to it the snappy way because this is ridiculous

OpenStudy (megannicole51):

hahah okay show me!

OpenStudy (anonymous):

do you have this \[\frac{x}{x^{2}-9x+20}=\frac{A}{x-4}+\frac{B}{x-5}\] somewhere on your paper?

OpenStudy (megannicole51):

yes

OpenStudy (megannicole51):

no

OpenStudy (megannicole51):

my 4 and 5 are switched but it doesnt really matter does it?

OpenStudy (anonymous):

oh ok is it switched? just want to make sure so we don't screw up

OpenStudy (megannicole51):

lol yeah bt keep goin!

OpenStudy (anonymous):

\[\frac{x}{x^{2}-9x+20}=\frac{A}{x-5}+\frac{B}{x-4}\]

OpenStudy (megannicole51):

:) good

OpenStudy (anonymous):

so also on your paper you have \[A(x-4)+B(x-5)=x\] right?

OpenStudy (megannicole51):

yup

OpenStudy (anonymous):

ok now here is the deal: \[A(x-4)+B(x-5)=x\] must be true NO MATTER WHAT \(x\) IS !

OpenStudy (anonymous):

in particalular it must be true if \(x=5\) and if \(x=5\) you get \[A(5-4)+B(5-5)=5\] i.e. \[A=5\]

OpenStudy (anonymous):

similarly it must be true if \(x=4\) which, skipping the stupid step of writing \(4-4\) we go right two \[B(4-5)=4\] or \(-B=4\) and so \(B=-4\) done finished no system etc

OpenStudy (megannicole51):

yeah thats how you do it when there isnt a coefficient after the equal sign

OpenStudy (megannicole51):

dang! i like it:)

OpenStudy (megannicole51):

thank you for showing me!

OpenStudy (anonymous):

and there is an even easier way, which amounts to the same thing

OpenStudy (megannicole51):

i love math lol

OpenStudy (anonymous):

\[\frac{x}{(x-5)(x-4)}=\frac{A}{x-5}+\frac{B}{x-4}\]

OpenStudy (megannicole51):

just find the zeros?

OpenStudy (anonymous):

find \(A\) the denominator of \(A\) is \(x-5\) so replace \(x\) by \(5\) \[\frac{x}{x-4}\]

OpenStudy (anonymous):

that is what i meant by this \[\frac{x}{\cancel{(x-5)}(x-4)}\] put your hand over the denominator of \(A\) and then replace \(x\) by \(5\)

OpenStudy (anonymous):

you get \(A=5\) in your head

OpenStudy (anonymous):

much much easier than multiplying out, collecting terms, equating coefficients, solving linear systems, it makes me exhausted just to think about it

OpenStudy (megannicole51):

lol agreed but thts wat i love about math

OpenStudy (anonymous):

glad you love it too btw the method i just showed obviously doesn't work all the time for example if you are going to end up with something like \[\frac{2x+1}{x^2+1}+\frac{x-2}{x^2+7}\] however you should always check to see if you can plug in anything for \(x\) to drop out all terms but one, making the solution to one of the constants simple

OpenStudy (megannicole51):

oh okay! good to know! thank you for your help!!!

OpenStudy (anonymous):

yw